Each vertex of the cube touches the edge of the sphere. If the cube is 2 m on each side, what is the radius of the sphere? Write your answer to two decimal places.
Answer by Alan3354(69443)

Each vertex of the cube touches the edge of the sphere. If the cube is 2 m on each side, what is the radius of the sphere? Write your answer to two decimal places.
---------
Diameter = sqrt(2^2 + 2^2 + 2^2) = sqrt(12) = 2sqrt(3)
radius = sqrt(3)